when both electric and magnetic field both are present , how can we findthe acceleration and postion of particle at a given time istant
 

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago
dear student
please use lorentz law and basic equations of kinematics like a=dv/dt,v=dx/dt,a=vdv/dx
hope u got it